OpenStudy (anonymous):

-2(y-4)-(3y-2)=-2-(6y-2)

OpenStudy (ivettef365):

First solve what is in parenthesis

OpenStudy (anonymous):

-2y-8-3y-2=-2-(6y-2)

OpenStudy (ivettef365):

actually it is -2y+8-3y+2 = -2-6y+2

OpenStudy (ivettef365):

Now combine like terms

OpenStudy (anonymous):

1y+6 = -2-(6y-2)?

OpenStudy (ivettef365):

-2y -3y = -5y 8 + 2 = 10 so you have -5y+10 = -2 +y +2 now do the same on the other side

OpenStudy (ivettef365):

sorry I meant -5y +10 = -2 -6y +2

OpenStudy (anonymous):

I do not see the +10 when -2y+(-3y)+8-2

OpenStudy (ivettef365):

lets start you have -2(y-4)-(3y-2) = -2-(6y-2) -2y+8-3y+4 = -2-6y+2

OpenStudy (ivettef365):

(-2)(y)=-2y (-2)(-4) = positive 8 (-)(3y) = -3y (-)(-2) = positive 2 can you see it ??

OpenStudy (ivettef365):

sorry it is -2y+8-3y+2 = -2-6y+2

OpenStudy (ivettef365):

now you combine like terms -2y-3y = -5y and 8+2 = 10

OpenStudy (ivettef365):

now you have -5y+10 = -6y (because -2+2=0)

OpenStudy (ivettef365):

Hope you can take it from there, I have to go

OpenStudy (anonymous):

Thank you for your help